Consider the linear array BBB[100] where lower bound is -10.
Suppose Base (BBB) = 300 and w = 4 words per memory cell for BBB. Find the
address of BBB [-10], BBB [0] and BBB [15].
1
Expert's answer
2020-12-13T22:56:25-0500

Using the formula: LOC(BBB[k]) = Base(BBB)+w(K-LB)

LOC(BBB[-10])= 300+4(-10-(-10)) = 300

LOC(BBB[0])= 300+4(0-(-10)) = 340

LOC(BBB[15[) = 300 + 4(15-(-10)) = 400